# Adaptive Off-Policy Inference for M-Estimators Under Model Misspecification

### v1

@@ 5 Empirical Results
We deploy these methods in the contextual bandit problem, where the goal is to choose a sequence of actions $\{A_{t}\}_{t=1}^{T}$ that minimize regret, $\sum_{t=1}^{T}\left(\mu^{\star}_{t}-Y_{t}\right)$, where $\mu^{\star}_{t}:=\max\{\mathbb{E}[Y_{t}(a)]:a\in\mathcal{A}\}$ is the arm with the highest expected reward. This is a well-studied problem with a variety of proposed solutions; we focus on strategies for choosing $P(A_{t}=a|X_{t},\mathcal{H}_{t-1})$ that simultaneously leverage flexible predictive models $f_{t}$ and $e_{t}$ described in 9 targetting $\mathbb{E}[Y_{t}|X_{t},A_{t}]$ and $\mathbb{E}[Y_{t}^{2}|X_{t},A_{t}]$ respectively.
Strategies for selecting $A_{t}$
@@
  • An upper confidence bound (UCB) strategy that constructs confidence intervals using $f_{t}$ to define the centerpoint and $v_{t}=e_{t}-f_{t}^{2}$ to control the width. Here,
    $A_{t}=\operatorname*{argmax}_{a\in\mathcal{A}}\left[f_{t}(X_{t},a)-q_{\alpha/2}\sqrt{v_{t}(X_{t},a)}\right],$
@@
  • An approach inspired by Thompson Sampling which draws from a distribution of $Y_{t}|X_{t},a$ for each $a$ to form an estimate $\hat{Y}_{t}(a)$ and then lets $A_{t}:=\operatorname*{argmax}\hat{Y}_{t}(a)$. We apply this algorithm in the current setting using the working assumption that $Y_{t}(a)\sim N(f_{t}(X_{t},a),e_{t}(X_{t},a)-f_{t}^{2}(X_{t},a))$. We also consider versions of this method with a clipping constraint so all probabilities are constrained to lie in $[0.05,0.95]$.
@@
We consider the outcome ($Y_{t}$) to be the four year change in KL grade for the affected knee of an osteoarthritic patient, as described in Dunn et al. (2020). The features ($X_{t}$) include the aforementioned baseline measurements of knee health and demographic variables. Since OAI was an observational study to characterize disease progression rather than a trial to compare treatments, to evaluate methodology, we construct a semi-synthetic data set that applies hypothetical treatments ($A_{t}$) and then enforces a synthetic relationship between $A_{t}$ and $Y_{t}$ while preserving the conditional distribution of $Y_{t}|X_{t}$. This is done using machine learning methods to first learn the distribution of $Y_{t}|X_{t}$ and then to create synthetic outcomes that combine the output of the machine learning model with hypothetical treatment effects. For a detailed description of the semi-synthetic dataset creation process, see Appendix C. The end result is a dataset with similar complexity to the original longitudinal study but with the option to sample patients sequentially with ground-truth knowledge of the treatment effect and target parameters. Note that to ensure that this setting would align with the hardest case where treatment policies would not be likely to converge, we picked treatments so that there would be the same expected reward across multiple arms and there would be no unique optimal policy.
(a) Coverage for target $1-\alpha=0.8$
@@
In simulations, patients come online one at a time and $\hat{\theta}_{t}$ is recomputed for each $t$ with Theorem 1 used to form confidence intervals. As discussed in Section 4, online model-based estimates of the conditional mean and variance are trained sequentially using random forests as inputs into the MAIPWMM estimator. To increase computational efficiency, we choose to re-train the predictive models after every $100$ data points, though in principle they could be re-trained after every sample. For continuous outcomes, we consider a linear model with $m_{\theta}:=(Y_{t}-\sum_{a\in\mathcal{A}}\theta_{a}\mathbbm{1}_{A_{t}=a})^{2}$. In the case of logistic regression, we threshold the outcome to a binary variable $\tilde{Y}_{t}:=\mathbbm{1}_{Y_{t}>\mathbb{E}\left[Y_{t}\right]}$ and choose $m_{\theta}:=\log(1+\exp(\sum_{a\in\mathcal{A}}\theta_{a}\mathbbm{1}_{A_{t}=a}))Y_{t}-(1-Y_{t})\log\left(1-\exp(\sum_{a\in\mathcal{A}}\theta_{a}\mathbbm{1}_{A_{t}=a})\right)$
We track the length of the confidence intervals and their nominal coverage rate (given target coverage of $0.8$) in Figure 2. The methods we consider are:
Methods to estimate $V_{t,\theta^{\star}}$
@@
  • MAIPWM–Naive Using the method described in Theorem 1 with plug-in estimates $V_{t,\theta^{\star}}$ found using the sample covariance of the score function (with no assumption of a time-varying component). That is, we let
    $\hat{V}_{t}:=\frac{1}{T}\sum_{t=1}^{T}s_{t,\hat{\theta}_{T}}s_{t,\hat{\theta}_{T}}^{T}.$
    This is a valid approach only when the variance is uniform across time (i.e. treatment propensities do not change across time) or a function of a converging policy.
  • IPW Use off-the-shelf maximum likelihood estimates with inverse propensity weighting and sandwich estimates of the variance.
We tested the MAIPWM estimator in all combinations of strategies used to estimate $V_{t,\theta^{\star}}$ and select actions. The empirical results shown in Figure 2 align with the theory — IPW estimates undercover in all situations other than the uniform (nonadaptive case). IPW as well as MAIPW estimates with naive estimates of the covariance require a significantly larger number of samples before they achieve the nominal coverage rate and do not cover at all in the case of Thompson sampling. The MAIPW using sample splitting and external data both cover correctly, with the sample splitting method paying a price in terms of notably wider confidence intervals. Although we have no theoretical results supporting this particular method, we also note that reusing the features from history to estimate the covariance appears to be indistinguishable from a method that uses external samples — future work can potentially justify this formally.
@@ 6 Conclusion
We present a method for inference in $M$ estimation problems under a potentially misspecified working model. The method requires the estimation of time-varying covariance matrices, for which we provide an algorithm for generalized linear models using flexible machine learning approaches. Empirical results demonstrate that the method had correct nominal coverage, while existing approaches often undercover in situations where action-selection probabilities vary considerably and the variance of the score function does not converge. Our work suggests several potential avenues for follow-up work, which we detail below:
@@
Throughout, we assume that the probabilities of each action being selected in the experiment is determined by the experimenter and known, which sometimes cannot be guaranteed in practical settings. Weakening this assumption first requires modifying the asymptotic arguments for Theorem 1 and ensuring the rate of estimation of these probabilities to converge sufficiently fast. It also complicates the estimation of the time-varying variance as explained in Section 4 since Proposition 1 crucially relies on knowledge of the action-selection probabilities to decompose the variance into a tractable form for plug-in estimation.
@@
Acknowledgments
Data and/or research tools used in the preparation of this manuscript were obtained and analyzed from the controlled access data sets distributed by the Osteoarthritis Initiative (OAI), a data repository located within the NIMH Data Archive (NDA). OAI is a collaborative informatics system created by the National Institute of Mental Health and the National Institute of Arthritis, Musculoskeletal and Skin Diseases (NIAMS) to provide a worldwide resource to quicken the pace of biomarker identification, scientific investigation and OA drug development. Dataset identifier(s): [NIMH Data Archive Collection ID(s) or NIMH Data Archive Digital Object Identifier (DOI)].
@@ Appendix A Preliminaries
We introduce several key facts about martingales that are used throughout. For a more thorough discussion, consult Hall and Heyde (2014)
@@
Let $\left\{S_{n}=\sum_{i=1}^{n}X_{i},\mathcal{H}_{n},n\geq 1\right\}$ be a martingale and $\left\{b_{n}\right\}$ a sequence of positive constants with $b_{n}\rightarrow\infty$ as $n\rightarrow\infty$. Then, writing $X_{ni}=X_{i}\mathbbm{1}\left[\left|X_{i}\right|\leq b_{n}\right]$, $1\leq i\leq n$, we have that $b_{n}^{-1}S_{n}\xrightarrow{\mathrm{p}}0$ as $n\rightarrow\infty$ if
@@
    $b_{n}^{-1}\sum_{i=1}^{n}\mathbb{E}\left[X_{ni}\mid\mathcal{H}_{t-1}\right]\xrightarrow{\mathrm{p}}0$
@@
    $b_{n}^{-2}\sum_{i=1}^{n}\left\{\mathbb{E}\left[X_{ni}^{2}\right]-\mathbb{E}\left[\mathbb{E}\left[X_{ni}\mid\mathcal{H}_{t-1}\right]\right]^{2}\right\}\rightarrow 0$
@@
Fact 2 (Martingale Central Limit Theore, Hall and Heyde (2014)).
Let $\left\{S_{ni},\mathcal{H}_{ni},1\leqslant i\leqslant k_{n},n\geqslant 1\right\}$ be a zero-mean, square integrable martingale array with differences $X_{ni}$, and let $\eta^{2}$ be an almost surely finite randomv ariable. Suppose that and
  1. 1.
    $\max_{i}\left|X_{ni}\right|\xrightarrow{\mathrm{p}}0$;
  2. 2.
    $\sum_{i}X_{ni}^{2}\xrightarrow{\mathrm{p}}\eta^{2}$;
  3. 3.
    $E\left(\max_{i}X_{ni}^{2}\right)$ is bounded in $n$.
Assume the $\sigma$-fields are nested: $\mathcal{H}_{n,i}\subseteq\mathcal{H}_{n+1,i}$ for $1\leq i\leq k_{n},\quad n\geq 1$. Then $S_{nk_{n}}=\sum_{i}X_{ni}\xrightarrow{\mathrm{~d}}N(0,1)$.
@@ B.1 Proof of Theorem 1
The overall structure of the proof is similar to Zhang et al. (2021), which in turn is is based on classical proofs of the asymptotic normality of $M$-estimators such as the ones described in van der Vaart (2000). The major differences in the approach are that we now need to deal with the variance stabilizing matrices $\hat{V}_{t}$ and bound the complexity of the predictive model $f_{t}$.
@@
    Lemma 3 show that $\left\lVert\hat{\theta}_{T}-\theta^{\star}\right\rVert_{1}=o_{p}(1)$
@@
$-\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}s_{t,\theta^{\star}}$ $=\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\left(-s_{t,\hat{\theta}_{T}}+\dot{s}_{t,\hat{\theta}_{T}}(\hat{\theta}_{T}-\theta^{\star})+\frac{1}{2}\left(\hat{\theta}_{T}-\theta^{\star}\right)^{T}\ddot{s}_{t,\tilde{\theta}_{t}}\left(\hat{\theta}_{T}-\theta^{\star}\right)\right).$
Recall that by assumption $\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}s_{t,\hat{\theta}_{T}}=0$, this simplifies the expression to
$-\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}s_{t,\theta^{\star}}$ $=\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\left(\dot{s}_{t,\hat{\theta}_{T}}(\hat{\theta}_{T}-\theta^{\star})+\frac{1}{2}\left(\hat{\theta}_{T}-\theta^{\star}\right)^{T}\ddot{s}_{t,\tilde{\theta}_{t}}\left(\hat{\theta}_{T}-\theta^{\star}\right)\right)$
$=\frac{1}{\sqrt{T}}\left(I_{d}+\frac{1}{2}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\left(\hat{\theta}_{T}-\theta^{\star}\right)^{T}\ddot{s}_{t,\tilde{\theta}_{t}}\left(\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\dot{s}_{t,\hat{\theta}_{T}}\right)^{-1}\right)\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\dot{s}_{t,\hat{\theta}_{T}}\left(\hat{\theta}_{T}-\theta^{\star}\right).$
@@
in order to conclude that $\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\dot{s}_{t,\hat{\theta}_{T}}\left(\hat{\theta}_{T}-\theta^{\star}\right)\overset{d}{\to}N(0,I_{d})$. We write this out as
@@
By Lemma 3, $\left\lVert\hat{\theta}_{T}-\theta^{\star}\right\rVert_{2}=o_{p}(1)$. Therefore, it is sufficient to show that the remaining terms are $O_{p}(1)$ to conclude that the entire term is $o_{p}(1)$. We know that $1/\delta_{\text{min}}$ is bounded because $\delta_{\text{min}}>0$ by assumption. We tackle each of these individually (note that by equivalence of matrix norms, it is sufficient to show convergence in $1$-norm).
@@
We have that
$\left\lVert\ddot{s}_{t,\tilde{\theta}_{t}}\right\rVert_{1}$ $=\left\lVert\sum_{a=1}^{K}\pi_{e}(A_{t}=a|X_{t})\left(\dddot{m}_{\theta}(a,X_{t},f_{t}(a,X_{t}))+\mathbbm{1}_{A_{t}=a}\frac{\dddot{m}_{\theta}(X_{t},A_{t},Y_{t})-\dddot{m}_{\theta}(X_{t},A_{t},f_{t}(X_{t},A_{t}))}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)\right\rVert$
$\leq\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}\left\lVert\dddot{m}_{\theta}(X_{t},A_{t},Y_{t})\right\rVert_{1}+\pi_{e}(A_{t}=a|X_{t})\sum_{a\in\mathcal{A}}\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)\left\lVert\dddot{m}_{\theta}(X_{t},a,f_{t}(X_{t},a))\right\rVert_{1}$
$\leq\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}u_{m}(X_{t},A_{t},Y_{t})+\pi_{e}(A_{t}=a|X_{t})\sum_{a\in\mathcal{A}}\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)u_{\mathcal{F}}(X_{t},a,f_{t}(X_{t},a)),$
where the last inequalities follow from 7 and 6. Therefore,
$\sum_{t=1}^{T}\left\lVert\ddot{s}_{t,\tilde{\theta}_{t}}\right\rVert\leq\sum_{t=1}^{T}\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}u_{m}(X_{t},A_{t},Y_{t})+\sum_{t=1}^{T}\pi_{e}(A_{t}=a|X_{t})\sum_{a\in\mathcal{A}}\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)u_{\mathcal{F}}(X_{t},a,f_{t}(X_{t},a).$
@@
$=C_{1}\mathbb{E}[\mathbb{E}_{\mathcal{P},p_{e}}[u_{m}(X_{t},A_{t},Y_{t})^{2}]]$
$=C_{1}\mathbb{E}_{\mathcal{P},p_{e}}[u_{m}(X_{t},A_{t},Y_{t})^{2}],$
@@
Similarly, we have $\mathbb{E}\left[\sum_{t=1}^{T}\pi_{e}(A_{t}=a|X_{t})\sum_{a\in\mathcal{A}}\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)u_{\mathcal{F}}(X_{t},a,f_{t}(X_{t},a)|\mathcal{H}_{t-1}\right]=0$. On the other hand, the variance can also be bounded. By 4 and the fact that $\pi_{e}(A_{t}=a|X_{t})<1$, we have that $\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)^{2}<(1-C_{2})^{2}$. Therefore,
@@
which is finite by assumption.
@@
Showing $\left\lVert\left(\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\dot{s}_{t,\hat{\theta}_{T}}\right)^{-1}\right\rVert_{2}=O_{p}(1)$.
It is sufficient to show that
$\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\dot{s}_{t,\hat{\theta}_{T}}\succeq H+o_{p}(1),$
for some positive definite matrix H. Expanding out the terms individually as before we have

### v2

@@ 5 Empirical Results
We deploy these methods in the contextual bandit problem, where the goal is to choose a sequence of actions $\{A_{t}\}_{t=1}^{T}$ that minimize regret, $\sum_{t=1}^{T}\left(\mu^{\star}_{t}-Y_{t}\right)$, where $\mu^{\star}_{t}:=\max\{\mathbb{E}[Y_{t}(a)]:a\in\mathcal{A}\}$ is the arm with the highest expected reward. This is a well-studied problem with a variety of proposed solutions; we focus on strategies for choosing $P(A_{t}=a|X_{t},\mathcal{H}_{t-1})$ that simultaneously leverage flexible predictive models $f_{t}$ and $j_{t}$ described in 9 targeting $\mathbb{E}[Y_{t}|X_{t},A_{t}]$ and $\mathbb{E}[Y_{t}^{2}|X_{t},A_{t}]$, respectively.
Strategies for selecting $A_{t}$:
@@
  • An upper confidence bound (UCB) strategy that constructs confidence intervals using $f_{t}$ to define the centerpoint and $j_{t}$ to control the width. Here,
    $A_{t}=\operatorname*{argmax}_{a\in\mathcal{A}}\left[f_{t}(X_{t},a)-q_{\alpha/2}\sqrt{j_{t}(X_{t},a)}\right],$
@@
  • An approach inspired by Thompson Sampling which draws from a distribution of $Y_{t}|X_{t},a$ for each $a$ to form an estimate $\hat{Y}_{t}(a)$ and then lets $A_{t}:=\operatorname*{argmax}\hat{Y}_{t}(a)$. We apply this algorithm in the current setting using the working assumption that $Y_{t}(a)\sim N(f_{t}(X_{t},a),j_{t}(X_{t},a)-f_{t}^{2}(X_{t},a))$. We also consider versions of this method with a clipping constraint so all probabilities are constrained to lie in $[0.05,0.95]$.
@@
We consider the outcome ($Y_{t}$) to be the four year change in KL grade for the affected knee of an osteoarthritic patient. Although KL grade is a discrete numeric variable, we treat KL grade as a continuous variable for the purposes of this example. The chosen features ($X_{t}$) include the aforementioned baseline measurements of knee health and demographic variables as well as additional risk factors identified in Dunn et al. (2020) such as self-reported quality of life scores and use of pain medication. Since OAI was an observational study to characterize disease progression rather than a trial to compare treatments, to evaluate methodology, we construct a semi-synthetic data set that applies hypothetical treatments ($A_{t}$) and then enforces a synthetic relationship between $A_{t}$ and $Y_{t}$ while preserving the conditional distribution of $Y_{t}|X_{t}$. This is done using machine learning methods to first learn the distribution of $Y_{t}|X_{t}$ and then to create synthetic outcomes that combine the output of the machine learning model with hypothetical treatment effects. For a detailed description of the semi-synthetic dataset creation process, see Appendix C. The end result is a dataset with similar complexity to the original longitudinal study but with the option to sample patients sequentially with ground-truth knowledge of the treatment effect and target parameters. Note that to ensure that this setting would align with the hardest case where treatment policies would not be likely to converge, we picked treatments so that there would be the same expected reward across multiple arms and there would be no unique optimal policy.
(a) Coverage for target $1-\alpha=0.9$
@@
In simulations, patients come online one at a time and $\hat{\theta}_{t}$ is recomputed for each $t$ with Theorem 1 used to form confidence intervals. As discussed in Section 4, online model-based estimates of the conditional mean and variance are trained sequentially (using random forests) as inputs into the MAIPWM estimator. To increase computational efficiency, we choose to re-train the predictive models after every $100$ data points, though in principle they could be re-trained after every sample. For continuous outcomes, we consider a linear model with $m_{\theta}:=-(Y_{t}-\sum_{a\in\mathcal{A}}\theta_{a}\mathbbm{1}_{A_{t}=a})^{2}$.
We track the length of the confidence intervals and their nominal coverage rate (given target coverage of $0.9$) in Figure 2. The methods we consider are:
Methods for confidence interval construction:
  • Naive Use off-the shelf maximum likelihood estimation with equal weighting of outcomes and sandwich estimates of the variance.
  • IPW Uses the methodology of Guo and Xu (2025) to construct confidence intervals, which relies on inverse propensity weighting of the score function and sandwich estimates of the variance.
  • SQ-IPW Use the methodology of Zhang et al. (2021) to construct confidence intervals, which weight observations by the square root of the propensity scores, setting $w_{t}=\sqrt{\frac{\pi_{e}(A_{t}|X_{t})}{P(A_{t}|X_{t},\mathcal{H}_{t-1})}}$ and using sandwich estimates of the variances.
@@
We tested the MAIPWM estimator in all combinations of strategies used to estimate $V_{t,\theta^{\star}}$ and select actions. The empirical results shown in Figure 2 align with the theory — naive maximum likelihood estimates undercover in all situations other than the uniform (nonadaptive case). IPW and SQ-IPW methods require a significantly larger number of samples before they achieve the nominal coverage rate and do not cover at all in the case of Thompson sampling. The MAIPW estimator using sample splitting and external data both cover correctly, with the sample splitting method paying a price in terms of notably wider confidence intervals. Although we have no theoretical results supporting MAIPWM-Reuse, which reuses the historical data to estimate the covariance, this method has nearly identical results to the two MAIPWM methods that use external samples. Future work can potentially justify this formally.
@@ 6 Conclusion
We present a method for inference in $M$ estimation problems under a potentially misspecified working model. The method requires the estimation of time-varying covariance matrices, for which we provide an algorithm for generalized linear models using flexible machine learning approaches. Empirical results demonstrate that the method has correct nominal coverage, while existing approaches often undercover in situations where action-selection probabilities vary considerably and the variance of the score function does not converge. Our work suggests several potential avenues for follow-up work, which we detail below:
@@
Throughout, we assume that the probabilities of each action being selected in the experiment are determined by the experimenter and known, which sometimes cannot be guaranteed in practical settings. Weakening this assumption first requires modifying the asymptotic arguments for Theorem 1 and ensuring the rate of estimation of these probabilities to converge sufficiently fast. It also complicates the estimation of the time-varying variance as explained in Section 4 since Proposition 1 crucially relies on knowledge of the action-selection probabilities to decompose the variance into a tractable form for plug-in estimation.
@@ Appendix A Preliminaries
We introduce several key facts about martingales that are used throughout. For a more thorough discussion, consult Hall and Heyde (2014).
@@
Let $\left\{S_{n}=\sum_{i=1}^{n}X_{i},\mathcal{H}_{n},n\geq 1\right\}$ be a martingale with respect to a filtration $\mathcal{H}_{n}$ and $\left\{b_{n}\right\}$ a sequence of positive constants with $b_{n}\rightarrow\infty$ as $n\rightarrow\infty$. Then, writing $X_{ni}=X_{i}\mathbbm{1}\left[\left|X_{i}\right|\leq b_{n}\right]$, $1\leq i\leq n$, we have that $b_{n}^{-1}S_{n}\xrightarrow{\mathrm{p}}0$ as $n\rightarrow\infty$ if
@@
    $b_{n}^{-1}\sum_{i=1}^{n}\mathbb{E}\left[X_{ni}\mid\mathcal{H}_{n-1}\right]\xrightarrow{\mathrm{p}}0$
@@
    $b_{n}^{-2}\sum_{i=1}^{n}\left\{\mathbb{E}\left[X_{ni}^{2}\right]-\mathbb{E}\left[\mathbb{E}\left[X_{ni}\mid\mathcal{H}_{n-1}\right]\right]^{2}\right\}\rightarrow 0$
@@ B.1 Proof of Theorem 1
The overall structure of the proof is similar to Zhang et al. (2021), which in turn is based on classical proofs of the asymptotic normality of $M$-estimators such as the ones described in van der Vaart (2000). The major differences in the approach are that we now need to deal with the variance stabilizing matrices $\hat{V}_{t}$ and bound the complexity of the predictive model $f_{t}$.
@@
    Lemma 4 shows that $\left\lVert\hat{\theta}_{T}-\theta^{\star}\right\rVert_{1}=o_{p}(1)$.
@@
$-\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}s_{t,\theta^{\star}}$ $=\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\left(s_{t,\hat{\theta}_{T}}-\dot{s}_{t,\hat{\theta}_{T}}(\hat{\theta}_{T}-\theta^{\star})-\frac{1}{2}\left(\hat{\theta}_{T}-\theta^{\star}\right)^{T}\ddot{s}_{t,\tilde{\theta}_{t}}\left(\hat{\theta}_{T}-\theta^{\star}\right)\right).$
Recall that by assumption $\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}s_{t,\hat{\theta}_{T}}=o_{p}(1/\sqrt{T})$, this simplifies the expression to
$-\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}s_{t,\theta^{\star}}$ $=o_{p}(1)+\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\left(\dot{s}_{t,\hat{\theta}_{T}}(\hat{\theta}_{T}-\theta^{\star})+\frac{1}{2}\left(\hat{\theta}_{T}-\theta^{\star}\right)^{T}\ddot{s}_{t,\tilde{\theta}_{t}}\left(\hat{\theta}_{T}-\theta^{\star}\right)\right)$
$=o_{p}(1)+\frac{1}{\sqrt{T}}\Bigg(I_{d}+\frac{1}{2}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}(\hat{\theta}_{T}-\theta^{\star})^{T}\ddot{s}_{t,\tilde{\theta}_{t}}$
$\qquad\qquad\qquad\qquad\times\Bigg(\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\dot{s}_{t,\hat{\theta}_{T}}\Bigg)^{-1}\Bigg)\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\dot{s}_{t,\hat{\theta}_{T}}(\hat{\theta}_{T}-\theta^{\star}).$
@@
in order to conclude that $\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\dot{s}_{t,\hat{\theta}_{T}}\left(\hat{\theta}_{T}-\theta^{\star}\right)\overset{d}{\to}N(0,I_{d})$. We write this as follows.
@@
By Lemma 4, $\left\lVert\hat{\theta}_{T}-\theta^{\star}\right\rVert_{2}=o_{p}(1)$ and by Lemma 2, $\left\lVert\left(\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\dot{s}_{t,\hat{\theta}_{T}}\right)^{-1}\right\rVert_{2}=O_{p}(1)$ . Therefore, it is sufficient to show that the remaining terms are $O_{p}(1)$ to conclude that the entire term is $o_{p}(1)$. We know that $1/\delta_{\text{min}}$ is bounded because $\delta_{\text{min}}>0$ by assumption. We tackle each of these terms individually (note that by equivalence of matrix norms, it is sufficient to show convergence in $1$-norm).
@@
We for some
$\left\lVert\ddot{s}_{t,\tilde{\theta}_{t}}\right\rVert_{1}$ $=\left\lVert\sum_{a=1}^{K}\pi_{e}(A_{t}=a|X_{t})\left(\dddot{m}_{\tilde{\theta}_{t}}(a,X_{t},f_{t}(a,X_{t}))+\mathbbm{1}_{A_{t}=a}\frac{\dddot{m}_{\tilde{\theta}_{t}}(X_{t},A_{t},Y_{t})-\dddot{m}_{\tilde{\theta}_{t}}(X_{t},A_{t},f_{t}(X_{t},A_{t}))}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)\right\rVert_{1}$
$\leq\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}\left\lVert\dddot{m}_{\tilde{\theta}_{t}}(X_{t},A_{t},Y_{t})\right\rVert_{1}+\pi_{e}(A_{t}=a|X_{t})\sum_{a=1}^{K}\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)\left\lVert\dddot{m}_{\tilde{\theta}_{t}}(X_{t},a,f_{t}(X_{t},a))\right\rVert_{1}$
$\leq\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}u_{m}(X_{t},A_{t},Y_{t})+\pi_{e}(A_{t}=a|X_{t})\sum_{a=1}^{K}\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)u_{\mathcal{F}}(X_{t},a,f_{t}(X_{t},a)),$
where the last inequalities follow from 7 and 6 after noting that $\left\lVert\tilde{\theta}_{t}-\theta^{\star}\right\rVert<\delta$ for any fixed $\delta$ since $\tilde{\theta}_{t}$ must lie on the line segment between $\hat{\theta}_{T}$ and $\theta^{\star}$. Therefore,
$\sum_{t=1}^{T}\left\lVert\ddot{s}_{t,\tilde{\theta}_{t}}\right\rVert$ $\leq\sum_{t=1}^{T}\frac{\pi_{e}(A_{t}\mid X_{t})}{\mathbb{P}\!\left(A_{t}\mid X_{t},\mathcal{H}_{t-1}\right)}\,u_{m}(X_{t},A_{t},Y_{t})$
$\quad+\sum_{t=1}^{T}\pi_{e}(A_{t}=a\mid X_{t})\sum_{a=1}^{K}\Bigg(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\!\left(A_{t}=a\mid X_{t},\mathcal{H}_{t-1}\right)}\Bigg)\,u_{\mathcal{F}}\!\left(X_{t},a,f_{t}(X_{t},a)\right).$
@@
$=C_{1}\mathbb{E}[\mathbb{E}_{\mathcal{P},\pi_{e}}[u_{m}(X_{t},A_{t},Y_{t})^{2}]]$
$=C_{1}\mathbb{E}_{\mathcal{P},\pi_{e}}[u_{m}(X_{t},A_{t},Y_{t})^{2}],$
@@
Similarly, we have
$\mathbb{E}\left[\sum_{t=1}^{T}\pi_{e}(A_{t}=a|X_{t})\sum_{a=1}^{K}\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)u_{\mathcal{F}}(X_{t},a,f_{t}(X_{t},a)|\mathcal{H}_{t-1}\right]$
$\leq\sum_{a=1}^{K}\mathbb{E}\left[\mathbb{E}\left[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)u_{\mathcal{F}}(X_{t},a,f_{t}(X_{t},a)|X_{t},\mathcal{H}_{t-1}\right]|,\mathcal{H}_{t-1}\right]$
$=\sum_{a=1}^{K}\mathbb{E}\left[\mathbb{E}\left[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)|X_{t},\mathcal{H}_{t-1}\right]u_{\mathcal{F}}(X_{t},a,f_{t}(X_{t},a)|\mathcal{H}_{t-1}\right]$
$=0,$
where the last line follows from the fact that $\mathbb{E}\left[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)|X_{t},\mathcal{H}_{t-1}\right]=0$.
On the other hand, the variance can also be bounded. By 4 and the fact that $\pi_{e}(A_{t}=a|X_{t})<1$, we have that $\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)^{2}<(1-C_{2})^{2}$. Therefore,
@@
which is finite by 6.
Lemma 2.
@@
Under the conditions of Theorem 1, $\left\lVert\left(\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}\dot{s}_{t,\hat{\theta}_{T}}\right)^{-1}\right\rVert_{2}=O_{p}(1)$.
Proof.
Let us decompose